OpenStudy (lacris):

Can anyone help me with this? I will give you a shiny medal :) Please give detailed instructions :> If Sinθ = 3/5 and 0≤ θ ≤π/2, then tan θ = to what? I'm not sure what to do so please give me detailed instructions, thank you :)

OpenStudy (lacris):

I do know that tan is opposite/adjacent? But I'm not sure how to deal with the inequality :,<

OpenStudy (agent0smith):

The inequality just tells you what quadrant the angle is in (you only need to know it, to know if tan θ is positive or negative) Use Sinθ = 3/5 to draw a right triangle, since sinθ = opposite/hyptenuse.

OpenStudy (lacris):



OpenStudy (lacris):

so how exactly do I find the adjacent length? It will be 3/x tan?

OpenStudy (lacris):

\[\frac{ 3 }{ x }\tan^{-1} \]

OpenStudy (agent0smith):

No. Use pythag.

OpenStudy (lacris):

ooh a^2+b^2 = c^2?

OpenStudy (lacris):

so its: \[3^2 + b^2 = 5^2\]

OpenStudy (lacris):

Which then becomes 9 + b^2 = 25 so I subtract 9 to 25 so I get B^2= 16

OpenStudy (lacris):

so x = 4 because the 4(4) = 16?

OpenStudy (lacris):

so it's tanθ = 3/4

OpenStudy (agent0smith):

Yep.

OpenStudy (lacris):

aww thank you so much! :D

OpenStudy (agent0smith):

But don't forget, which quadrant is this? 0≤ θ ≤π/2 And is tan positive or negative in that quadrant?

OpenStudy (lacris):

I'm not really sure to be honest

OpenStudy (lacris):

it would be positive right? Im guessing? because:

OpenStudy (lacris):

Is that correct or no?

OpenStudy (agent0smith):

Not quite. This shows the angles (make sure you know this) http://thepythongamebook.com/_media/en:part2:600px-unit_circle_angles.svg-kleiner.png This show whether positive or negative http://cdn-3.ask-math.com/images/astc-rule-in-trigonometry.png

OpenStudy (lacris):

Ohh, I really screwed up on my drawing then oops so it would be positive because tan and cot are positive right? :D

OpenStudy (agent0smith):

Which quadrant is this? 0 ≤ θ ≤ π/2

OpenStudy (lacris):

and since its less than or equal to pi/2 that means its positive as well :D?

OpenStudy (lacris):

that quadrant is quadrant I right?

OpenStudy (agent0smith):

Yep

OpenStudy (lacris):

wow!! Omigosh, thank you so much for teaching me and helping me with this problem! I learned a lot!

OpenStudy (lacris):

My teachers didn't help me like that xD q-q

OpenStudy (agent0smith):

Welcome. Don't forget that they're all positive in quadrant 1. Probably cos not all math teachers really know math ;P

OpenStudy (agent0smith):

Welcome :)

OpenStudy (lacris):

:D

OpenStudy (agent0smith):

:D